OpenStudy (anonymous):

complete the graph of values for the function f(x)=3^x and sketch a graph of f(x) x=-2 -1 0 1 2

OpenStudy (anonymous):



OpenStudy (anonymous):

Similarly find f(x) when x = -1, 0, 1, 2

OpenStudy (anonymous):

thank youso lets just say f(-1)3^-1

OpenStudy (anonymous):

Yes, f(-1) = 3^-1 = 1/3

OpenStudy (anonymous):

tHx my ansers are -2 =(2,-9) , -1= (1,-3) , 0 =(0,-1) , 1 = (1,3), 2 =(2,9)